Carnations cost $23.75 per dozen, and roses cost $69.95 per dozen. The florist sold a combination of 12 dozen flowers on Saturday and took in $608.40. How many dozens of each kind of flower did the florist sell?
Answer by checkley77(12569)

69.95x+23.75(12-x)=608.40
69.95x+285-23.75x=608.40
46.2x=608.40-285
46.2x=232.4
x=232.4/46.2
x=7 doz. roses were sold.
12-7=5 doz. carnations were sold.
Proof:
69.95*7+23.75*5=608.40
489.65+118.75=608.40
6-8.40=608.40